OpenStudy (anonymous):

How would you solve this? n - 4 - 9

OpenStudy (freckles):

solve or simplify?

OpenStudy (freckles):

solve means we have an equation or inequality

OpenStudy (freckles):

we do not have that here

OpenStudy (freckles):

-4-9=-(4+9)=-13 so n-4-9=n-13

OpenStudy (♪chibiterasu):

Yeah I would assume simplification . . . so just add the regular numbers like freckles did. There's only one variable. Any further simplification would require a given solution.

OpenStudy (freckles):

I have to ask @beckyg111 did you mean n-4=9

OpenStudy (freckles):

or n-4=-9

OpenStudy (freckles):

or did you really just mean n-4-9

OpenStudy (anonymous):

ok in class im doing like the rainbow thin. Do you know what that is?

OpenStudy (anonymous):

It does say n - 4 - 9

OpenStudy (freckles):

ok in the instructions is to just simplify right?

OpenStudy (freckles):

and no never heard of rainbow thin sounds like the rainbow is sick :p

OpenStudy (anonymous):

heres another example 

OpenStudy (freckles):

i'm not familiar with that kind of mathematics

OpenStudy (freckles):

is that suppose to be -3x+6=-3x+2+4 ?

OpenStudy (♪chibiterasu):

are you sure its not like n(-4 - 9)?

OpenStudy (anonymous):

no. you bring the -3 and the x down so that would look like this. -3x. The 3 is negative and the 6 is positive so the result will be negative...i think then all you do is multiply 6 and 3 (my mistake its not 24 its 18) So then you bring the -3x and add the 18 -3x 18

OpenStudy (freckles):

so you mean -3(x+6) which when using distributive property gives -3(x)+(-3)(6) -3x+(-18) -3x-18

OpenStudy (anonymous):

I guess but i have to show my work and i have to do it the way my teacher showed me http://easyk12.blogspot.com/2011/10/multiplication-rainbow-method.html

OpenStudy (♪chibiterasu):

Yeah . . . The solution -3x - 18 does match with that.

OpenStudy (freckles):

omg

OpenStudy (freckles):

wait a minute

OpenStudy (freckles):

nevermind I'm confused again

OpenStudy (freckles):

like looking at that site almost thought x was a multiplication symbol here

OpenStudy (freckles):

instead of a variable

OpenStudy (♪chibiterasu):

Your teacher uses the most unnecessary methods . . .

OpenStudy (♪chibiterasu):

Ok I kinda get it now but I still have no clue . . .

OpenStudy (♪chibiterasu):

I didn't think I could struggle with this kind of math and I'm already in advanced math.

OpenStudy (♪chibiterasu):

She's gone lol.

OpenStudy (freckles):

I honestly don't think it is you. I think something is missing... because the site becky linked us too shows something completely different than problems expressed here.

OpenStudy (freckles):

3x(-6) where x is a multiplication symbol 3x(-6)=-18

OpenStudy (freckles):

I don't see what this has to do with n-4-9

OpenStudy (freckles):

actually she wrote -3x(+6) which is just really -3x(6) again x is a multiplication symbol this is still -18 though

OpenStudy (♪chibiterasu):

Yeah I have no idea why the "rainbow method" would apply to this.

OpenStudy (♪chibiterasu):

I feel like her teacher just made it even more complicated.